OpenStudy (anonymous):

Someone, please help? Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (johnweldon1993):

Well okay lets choose an easy one... Standard form 3x + 3y = 6 How would you find a point on this line?

OpenStudy (anonymous):

You would turn it in to slope form... right??

OpenStudy (johnweldon1993):

Well that's what they want you to do for part D... But yeah you can do that...so what would that be in slope intercept form?

OpenStudy (anonymous):

3y = -3x +6

OpenStudy (johnweldon1993):

Close...1 more step...remember slope intercept form is y = mx + b...notice how you have 3y....so to solve for 'y' you would...?

OpenStudy (anonymous):

I think you would divide it all by 3..

OpenStudy (johnweldon1993):

Right...which would then be...?

OpenStudy (anonymous):

y = -3/3x + 2

OpenStudy (johnweldon1993):

Right...or easier y = -x + 2 right?

OpenStudy (anonymous):

But I thought you kept the x a fraction for rise/run?

OpenStudy (johnweldon1993):

You can if you need to see it that way...so it would be y = -1/1x + 2 right? because -3/3 simplifies down to that... so if you need to see the rise/run ...then absolutely keep it like that...

OpenStudy (anonymous):

I don't have to, I just thought that's how you were suppose to.. Either way I'm following!

OpenStudy (johnweldon1993):

Alright cool, y = -x + 2 so part A is done..that was the standard form.....3x + 3y = 6 Part B, find a point on this line...well for that all you need to do is choose any random number for 'x' and then solve for 'y'...so say we have 3(3) + 3y = 6 9 + 3y = 6 solve that for 'y'...

OpenStudy (anonymous):

Wouldn't it be 3y= -9 + 6 then be y = -3 + 2 ??

OpenStudy (johnweldon1993):

continue on...what is -3 + 2 ?

OpenStudy (anonymous):

Oh.. -1?

OpenStudy (johnweldon1993):

Right! So when I made x = 3....then you solved and figured out that y = -1..so a point on this line would be (3 , -1) And there's part B ...okay?

OpenStudy (anonymous):

Okay, so part B is just (3, -1) ?

OpenStudy (johnweldon1993):

Right...and all the work we just did...since it does say justify the answer...

OpenStudy (anonymous):

Oh, right can't forget that

OpenStudy (johnweldon1993):

So part C now....to find the 'x' and 'y' intercepts...you need to find what the equation would equal when either 'x' or 'y' = 0 So for example....when the graph passes through the x-axis...this is when y = 0....so in our equation... 3x + 3y = 6 when y = 0 we have 3x + 3(0) = 6 3x = 6 solve for 'x' ....?

OpenStudy (anonymous):

Okay, that would be.... 2

OpenStudy (johnweldon1993):

Right...so the x-intercept would be at (2 , 0) And to solve for the y-intercept...the 'x' would = 0.. so 3x + 3y = 6 would become 3(0) + 3y = 6 3y = 6 and solve for 'y'...

OpenStudy (anonymous):

that would also be 2

OpenStudy (johnweldon1993):

Correct...so together we have x-intercept = (2 , 0) y-intercept = (0 , 2) Make sense?

OpenStudy (anonymous):

Yes, it does!

OpenStudy (johnweldon1993):

And finally Part D we already did.... y = -x + 2 right?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

to graph...we already know the x and y intercepts...so we would plot those 2 points and then connect them...  And that would be our graph

OpenStudy (anonymous):

Okay thank you so much! Could you help me with one more please?

OpenStudy (johnweldon1993):

I actually have to head out for about an hour or so...but go ahead and post it up and I'm sure someone else can help you out...if not...feel free to message me and I'll help you out when I get back home...sound good?

OpenStudy (anonymous):

That sounds good, thanks!

OpenStudy (johnweldon1993):

Anytime :)